# From Frege to chatGPT: Compositionality in language, cognition, and deep neural networks

### v1

@@ 5.2 Implications for Human Development
Metalearning The metalearning approach, on the other hand, relaxes the assumption that these inductive biases need to be characterizable in architectural terms at all, and explores how they might themselves be learned from data. As has been articulated clearly in other work (Binz et al.,, 2023; Lake and Baroni,, 2023; McCoy et al.,, 2020; Wang,, 2020), the metalearning approach is agnostic about whether this occurs on an evolutionary timescale or within an individual human lifetime. Many have shied away from interpreting metalearning neural networks as models of human development (although see Russin et al., 2024a, ; Wang,, 2020; Wang et al.,, 2018), instead choosing to emphasize its utility as a methodological tool (Binz et al.,, 2023; McCoy et al.,, 2020).
@@
A more specific prediction of the metalearning hypothesis is that inner-loop, in-context learning, as opposed to outer-loop, in-weight learning, should improve with experience. This is also consistent with existing evidence, which suggests that older children are much more adept at learning and reasoning in context using working memory and other executive functions (Munakata et al.,, 2012). Indeed, executive functions, which are intimately related to our ability to learn in context (Cole et al.,, 2011; Duncan et al.,, 2017; Hampshire et al.,, 2011; Miller and Cohen,, 2001; Rougier et al.,, 2005), emerge especially late in development and continue to mature throughout adolescence and early adulthood (Ferguson et al.,, 2021).
@@
It is important to note that humans learn from experiences that are highly unstructured compared to the training datasets used in most studies on metalearning (e.g., Lake,, 2019; Lake and Baroni,, 2023; Russin et al., 2024b, ), which are contrived by researchers to engender very specific inductive biases on very specific tasks. While humans are trained on structured curricula in formal educational settings, there is evidence that competences related to compositionality such as sensitivity to hierarchical structure in syntax (Chomsky,, 1957) and to geometric features in visual tasks (Dehaene et al.,, 2022) do not depend on such experiences. However, LLMs have shown that simply predicting the next word on a large dataset of unstructured text can engender an inductive bias for compositionality (Brown et al.,, 2020; Bubeck et al.,, 2023; Zhou et al., 2023a, ). These models develop in-context learning algorithms (von Oswald et al.,, 2023; Xie et al.,, 2022) that demonstrate good generalization performance on compositional tasks (Webb et al.,, 2023; Wei et al.,, 2023; Zhou et al., 2023a, ), even though their training sets were not specifically designed to achieve this. These results are impressive, and provide a proof of concept that an inductive bias for compositionality can be learned from unstructured language data.
However, current LLMs are trained on orders of magnitude more data than humans experience in an entire lifetime (Frank,, 2023; Linzen and Baroni,, 2021; Warstadt and Bowman,, 2024), making it unclear whether similar capabilities could emerge in models trained on a more realistic scale. The emergence of in-context learning abilities has been shown to depend on the distributional properties of the training data and on network architecture (Chan et al.,, 2022). Furthermore, architectural inductive biases such as those related to the “relational bottleneck” (Webb et al.,, 2024) have been shown to improve sample efficiency on certain tasks (Webb et al.,, 2021). Thus, perhaps human compositionality is best characterized as an emergent property of an inner-loop, in-context learning algorithm, but the brain embodies important architectural inductive biases that encourage this algorithm to emerge.
@@ 5.3 Mere Implementations?
The empirical evidence reviewed up to this point suggests that neural network models may be able to get past the first horn of Fodor and Pylyshyn’s critical dilemma (see section 2.3). We have focused on this first horn because we believe that recent advances represent a crucial development on this front, eliminating a major obstacle that has hindered the acceptance of neural networks as viable models of human cognition. But what about the ‘mere implementation’ objection, the second horn of the compositionality challenge? Is the success of contemporary DNNs on compositional generalization problems a reflection of the fact that they have implemented classical architectures? And if so, does this mean that they are uninformative for cognitive theory?
@@
Though this is a nuanced and challenging issue, certain cases are unambiguous. As mentioned in section 4.2, neuro-symbolic hybrid models are ‘mere implementations’ by design—proponents of this approach treat the behavioral signatures of compositionality in humans as diagnostic of underlying symbolic mechanisms and set out to explicitly build such mechanisms into neural networks. Though the neural components of these networks are utilized to deal with high-dimensional inputs like images, their compositionality is entirely due to their built-in symbolic components. [^23] This reflects a corresponding hypothesis about human cognition—that it is a combination of two distinct systems, one symbolic and one statistical—which has played a key role in, e.g., the past tense debate (Kirov and Cotterell,, 2018; Pinker and Prince,, 1988; Pinker and Ullman,, 2002). This captures something of the spirit of Fodor and Pylyshyn’s original objection (in fact, they allude to this possibility themselves), and turns it into an empirical hypothesis with a clear predictive upshot—explicit implementation of symbolic mechanisms will be a necessary prerequisite for neural networks to succeed on compositional problems. Current results do not seem to favor this hypothesis, as the findings on strictly neural architectures (section 4.2), metalearning (section 4.3), and LLMs (section 4.4) suggest that such explicit symbolic mechanisms may not be necessary. However, it is still unclear whether neuro-symbolic hybrids will win out in the end. If so, it would mean that DNNs yield few new insights into compositionality, though they may still offer alternative perspectives on other aspects of cognition. [^24] Chalmers, (1993) makes this point. Though compositionality is an important explanatory target, it is just one of the phenomena that an adequate theory of mind must account for. He draws an extended analogy between Fodor and Pylyshyn’s dilemma and a behaviorist who argues that, because a cognitivist will need to account for classical conditioning along behaviorist lines, their approach is not a genuine alternative to the mind and cannot yield new insights. In both cases, the point is the same: local agreement about specific cases need not undermine the robust differences between two explanatory paradigms.
@@
However, the possibility that neural networks are implementing classical symbolic operations is invoked to stave off challenges from DNNs, even in the absence of any tangible predictions of this kind. Without clear mechanistic commitments, it becomes harder to assess. The results reviewed here (sections 4.2, 4.3, and 4.4) indicate that even strictly neural models with no built-in symbolic components may be adequate for replicating the behavioral signatures of compositionality. What should we make of this outcome?
@@
One possibility that was widely discussed in the original connectionist debate is eliminativism (Ramsey et al.,, 1990). An eliminative connectionist would argue that the success of strictly neural models demonstrates that Chomsky and Montague’s shift to treating formal syntactic and semantic theories as models of human psychological processing (see section 2) was a grave mistake (Piantadosi,, 2023), and that the ideal language philosophers’ skepticism about this approach was well-founded. If DNNs can achieve compositional generalization without internally representing any rule-based, combinatorial systems, this undermines the motivation for positing them to explain human behavior. Chomsky was right that this capacity cries out for explanation, but the eliminativist holds that the neural network vocabulary of nodes, weights, layers, attention heads, and so forth, will furnish sufficient explanatory resources for doing so.
@@
The possibility that these networks are implementing classical architectures serves as an obstacle to such eliminativist conclusions (McGrath et al., 2023a, ; McGrath and Russin,, 2024). Physical systems can exhibit functional organization at different levels of description. The fact that no explicit symbols or syntactic operations were built into a transformer that has metalearned to generalize compositionally doesn’t mean that it hasn’t learned to implicitly implement these classical constructs through training. For all we know—runs this version of the objection—the transformers that succeed on these compositional tasks are not alternative cognitive models at all, but mere implementations of the very same theoretical constructs that classical theorists have emphasized all along. On this view, they may help us to understand how classical architectures are implemented in the human brain (McLaughlin,, 1993), but will not be informative at the cognitive level.
@@
The only way to guarantee that the success of DNNs reflects mere implementation would be to treat the behavioral signatures of compositionality as constitutive, rather than diagnostic, of an underlying classical architecture (McGrath et al., 2023b, ). What it is to be a classical architecture, on this constitutive view, just is to be the kind of system that generates this behavioral profile. Adopting this position would ensure that any successful DNN would count as a mere implementation of a classical architecture, but it would do so at the cost of sapping classical theory of the kind of explanatory value that it was originally intended to contribute. The rise of cognitivism and its attendant rejection of behaviorism hinged on the claim that psychological theory must do more than merely redescribe behavior—we need an understanding of the independently identifiable underlying mechanisms that produce the given behavior (Cao and Yamins, 2021a, ). If giving rise to these generalization behaviors is constitutive of being a classical architecture, however, then we preclude even the logical possibility that a DNN could replicate the behavioral signatures of compositionality without implementing a classical system. This means, in effect, that we have failed to specify any theoretical content beyond this characterization of behavior and have lapsed back into mere redescription.
@@
Rather than simply assuming that contemporary DNNs implement classical architectures or searching for some a priori guarantee, what we need to do is try to find out whether or not they are doing so. The burgeoning field of “mechanistic interpretability” is making this increasingly feasible (Elhage et al.,, 2021; Olsson et al.,, 2022; Merullo et al.,, 2024; Wu et al.,, 2023). Though interpretability work is often raised in the context of AI safety considerations (Rudner and Toner,, 2021), it can also play a crucial role in informing cognitive theory, allowing us to extract higher-level, functionally relevant descriptions of these models’ internal representations that we can put into contact with existing cognitive theory (McGrath et al., 2023b, ; Millière,, 2024; Millière and Buckner, 2024b, ). In principle, this process could point us towards eliminativism, if successful models’ representations bear no resemblance to classical constructs, or mere implementationism, if they exactly recapitulate the structure of a fully specified symbolic theory. In the cases where this work has been pursued, however, it seems likely that these models land somewhere in between, recapitulating certain aspects of classical theory, but diverging in interesting ways (Baroni,, 2022; Chen et al.,, 2024; Manning et al.,, 2020; McGrath et al., 2023a, ). In such cases, we have suggested that it may be worth preserving key terms to encompass these behaviors (McGrath et al., 2023a, ; McGrath and Russin,, 2024), as long as it does not lose track of the fact that they contain new insights. In a slogan, not all systems that we may count as implementations are ‘mere’—informative implementations can point us to new insights into previously postulated structures (Smolensky,, 1988; Smolensky and Legendre,, 2011). Though interpretive work specifically focused on evaluating the compositional models reviewed here is in its initial stages, it is a promising line of future investigation (Andreas,, 2019; Lepori et al.,, 2023; Lewis et al.,, 2023; McCoy et al.,, 2019; Russin et al., 2021a, ; Soulos et al.,, 2020; Todd et al.,, 2024). From a philosophical perspective, further progress on deeper issues about the nature of explanation and implementation will put us in a position to use these findings to resolve the mere implementation objection once and for all.
@@ 5.4 Conclusion
In this article, we have traced the notion of compositionality from its emergence in early analytic philosophy to its present status as a key explanatory construct in contemporary cognitive science. We have emphasized the role that compositionality plays in debates over the explanatory adequacy of neural network models of human cognition and reviewed a range of recent results suggesting that current deep neural networks are capable of replicating some of the behavioral signatures that have motivated the ascription of internal compositional systems to human beings. In particular, these recent results suggest that architectural inductive biases and metalearning (or large-scale pretraining) are viable approaches to reproducing these behavioral signatures with strictly neural architectures. We have discussed how, when taken seriously as approaches to modeling human compositionality, these frameworks generate concrete predictions about its underlying neural mechanisms and about its development. This recent success also reopens key philosophical questions about the theoretical significance and proper interpretation of neural networks—whether they should be understood as merely implementing classical compositional mechanisms, or may generate novel insights and motivate revisions to classical theories. These pressing questions take us beyond asking what neural networks can do—instead asking what they mean for our understanding of the human mind. We expect that resolving them will require an interdisciplinary approach and increased philosophical engagement.
@@ 5.5 Acknowledgments
We would like to thank Richard Kimberley Heck, Josh Schechter, Christopher Hill, Elizabeth Miller, Roman Feiman, Ellie Pavlick, Michael J. Frank, Randall O’Reilly, Edouard Machery, Suraj Anand, Steven Frankland, Taylor Webb, Stavros Orfeas Zormpalas, and the participants in Brown University Dissertation Workshop and the Philosophy and Cognitive Science of Deep Learning reading group for helpful comments and discussions. JR was supported by NIH NIGMS COBRE grant GR5271961.

### v2

@@ 4.2 Metalearning and Human Development
The metalearning approach emphasizes how inductive biases can themselves be learned from data. As has been articulated clearly in other work (Binz et al.,, 2023; Lake and Baroni,, 2023; McCoy et al.,, 2020; Wang,, 2020), the metalearning approach is agnostic about whether this occurs on an evolutionary timescale or within an individual human lifetime. Many have shied away from interpreting metalearning neural networks as models of human development (although see Russin et al., 2024a, ; Wang,, 2020; Wang et al.,, 2018), instead choosing to emphasize its utility as a methodological tool (Binz et al.,, 2023; McCoy et al.,, 2020).
@@
A more specific prediction of the metalearning hypothesis is that ICL, as opposed to IWL, should improve with experience. This is also consistent with existing evidence, which suggests that older children are much more adept at learning and reasoning in context using working memory and other executive functions (Munakata et al.,, 2012). Indeed, executive functions, which are intimately related to our ability to learn in context (Cole et al.,, 2011; Duncan et al.,, 2017; Hampshire et al.,, 2011; Miller and Cohen,, 2001; Rougier et al.,, 2005), emerge especially late in development and continue to mature throughout adolescence and early adulthood (Ferguson et al.,, 2021).
@@
It is important to note that humans learn from experiences that are highly unstructured compared to the training datasets used in most studies on metalearning (e.g., Lake,, 2019; Lake and Baroni,, 2023; Russin et al., 2024b, ), which are contrived by researchers to engender very specific inductive biases on very specific tasks. While humans are trained on structured curricula in formal educational settings, there is evidence that competences related to compositionality such as sensitivity to hierarchical structure in syntax (Chomsky,, 1957) and to geometric features in visual tasks (Dehaene et al.,, 2022) do not depend on such experiences. However, LLMs have shown that simply predicting the next word on a large dataset of unstructured text can engender an inductive bias for compositionality (Brown et al.,, 2020; Bubeck et al.,, 2023; Zhou et al., 2023a, ). These models develop ICL abilities (von Oswald et al.,, 2023; Xie et al.,, 2022) that allow good generalization performance on compositional tasks (Webb et al.,, 2023; Wei et al.,, 2023; Zhou et al., 2023a, ), even though their training sets were not specifically designed to achieve this. These results are impressive, and provide a proof of concept that an inductive bias for compositionality can be learned from unstructured language data.
However, current LLMs are trained on orders of magnitude more data than humans experience in an entire lifetime (Frank,, 2023; Linzen and Baroni,, 2021; Warstadt and Bowman,, 2024), making it unclear whether similar capabilities could emerge in models trained on a more realistic scale. The emergence of ICL abilities has been shown to depend on the distributional properties of the training data and on network architecture (Chan et al.,, 2022). Furthermore, architectural inductive biases such as those related to the “relational bottleneck” (Webb et al.,, 2024) have been shown to improve sample efficiency on certain tasks (Webb et al.,, 2021). Thus, perhaps human compositionality is best characterized as an emergent property of an in-context learning algorithm, but the brain embodies important architectural inductive biases that encourage this algorithm to emerge.
@@ 4.3 Mere Implementations?
The empirical evidence reviewed up to this point suggests that neural network models may be able to get past the first horn of Fodor and Pylyshyn’s critical dilemma (see section 2.3). We have focused on this initial empirical challenge because we believe that recent advances represent a crucial development on this front, eliminating a major obstacle that has hindered the acceptance of neural networks as viable models of human cognition. But what about the ‘mere implementation’ objection, the second horn of the compositionality challenge? Is the success of contemporary DNNs on compositional generalization problems a reflection of the fact that they have implemented classical architectures? And if so, does this mean that they are uninformative for cognitive theory?
@@
Though this is a nuanced and challenging issue, certain cases are unambiguous. Neuro-symbolic hybrid models (Ellis et al.,, 2021; Gaur et al.,, 2022; Marcus,, 2020; Marcus and Davis,, 2020; Yu et al.,, 2023; Zhou et al.,, 2024), which combine explicit symbolic processes with neural networks, are ‘mere implementations’ by design—proponents of this approach treat the behavioral signatures of compositionality in humans as diagnostic of underlying symbolic mechanisms and explicitly build such mechanisms into neural networks. Though the neural components of these networks are utilized to deal with high-dimensional inputs like images, their compositionality is entirely due to their built-in symbolic components. This captures something of the spirit of Fodor and Pylyshyn’s original objection (in fact, they allude to this possibility themselves), and transforms it into an empirical hypothesis with clear predictive upshot—explicit implementation of symbolic mechanisms will be a necessary prerequisite for neural networks to succeed on compositional problems. Current results do not seem to favor this hypothesis, as the findings on metalearning (section 3.2), and LLMs (section 3.3) suggest that such explicit symbolic mechanisms may not be necessary. However, it is still unclear whether neuro-symbolic hybrids will win out in the end. If so, it would mean that DNNs yield few novel insights into compositionality, though they may still offer alternative perspectives on other aspects of cognition (Chalmers,, 1993).
@@
However, the claim that neural networks may implement classical symbolic operations is invoked to stave off challenges from DNNs, even in the absence of any tangible predictions of this kind. Without clear mechanistic commitments, it becomes harder to assess. The results reviewed here indicate that even strictly neural models with no built-in symbolic components may be adequate for replicating the behavioral signatures of compositionality. What should we make of this outcome?
@@
One possibility that was widely discussed in the original connectionist debate is eliminativism (Ramsey et al.,, 1990). An eliminative connectionist would argue that the success of strictly neural models demonstrates that Chomsky and Montague’s shift to treating formal syntactic and semantic theories as models of human psychological processing (see section 2) was a grave mistake (Piantadosi,, 2023), and that the ideal language philosophers’ skepticism about this approach was well-founded. If DNNs can achieve compositional generalization without internally representing any rule-based, combinatorial systems, this undermines the motivation for positing them to explain human behavior. Chomsky was right that this capacity cries out for explanation, but the eliminativist holds that the neural network vocabulary of nodes, weights, layers, and so forth, will furnish sufficient explanatory resources for doing so.
@@
The possibility that these networks are implementing classical architectures serves as an obstacle to such eliminativist conclusions (McGrath et al., 2023a, ; McGrath and Russin,, 2024). Physical systems can exhibit functional organization at different levels of description. The fact that no explicit symbols or syntactic operations were built into a DNN that has metalearned to generalize compositionally doesn’t mean that it hasn’t learned to implicitly implement these classical constructs through training. For all we know—runs this version of the objection—the DNNs that succeed on these compositional tasks are not alternative cognitive models at all, but mere implementations of the very same theoretical constructs that classical theorists have emphasized all along. On this view, they may help us to understand how classical architectures are implemented in the human brain (McLaughlin,, 1993), but will not be informative at the cognitive level.
@@
The only way to guarantee that the success of DNNs reflects mere implementation would be to treat the behavioral signatures of compositionality as constitutive (rather than diagnostic) of an underlying classical architecture (McGrath et al., 2023b, ). What it is to be a classical architecture, on this constitutive view, just is to be the kind of system that generates this behavioral profile. Adopting this position would ensure that any successful DNN would count as a mere implementation of a classical architecture, but it would do so at the cost of sapping classical theory of the kind of explanatory value that it was originally intended to contribute. The rise of cognitivism and its attendant rejection of behaviorism hinge on the claim that psychological theory must do more than merely redescribe behavior—we need an understanding of the independently identifiable underlying mechanisms that produce the given behavior (Cao and Yamins, 2021a, ). If giving rise to these generalization behaviors is constitutive of being a classical architecture, however, then we preclude even the logical possibility that a DNN could replicate the behavioral signatures of compositionality without implementing a classical system. This means, in effect, that we have failed to specify any theoretical content beyond this characterization of behavior and have lapsed back into mere redescription.
@@
Rather than simply assuming that contemporary DNNs implement classical architectures or searching for some a priori guarantee, what we need to do is try to find out whether or not they are doing so. The burgeoning field of “mechanistic interpretability” is making this increasingly feasible (Elhage et al.,, 2021; Olsson et al.,, 2022; Merullo et al.,, 2024; Wu et al.,, 2023). Though interpretability work is often raised in the context of AI safety considerations (Rudner and Toner,, 2021), it can also play a crucial role in informing cognitive theory, allowing us to extract high-level, functionally relevant descriptions of these models’ internal representations that we can put into contact with existing cognitive theory (McGrath et al., 2023b, ; Millière,, 2024; Millière and Buckner, 2024b, ). In principle, this process could point us towards eliminativism (if successful models’ representations bear no resemblance to classical constructs), or mere implementationism (if they exactly recapitulate the structure of a fully specified symbolic theory). In the cases where this work has been pursued, however, it seems likely that these models land somewhere in between, recapitulating certain aspects of classical theory, but diverging in interesting ways (Baroni,, 2022; Chen et al.,, 2024; Manning et al.,, 2020). In such cases, we have suggested that it may be worth preserving key terms to encompass these behaviors (McGrath et al., 2023a, ; McGrath and Russin,, 2024), as long as it does not obscure the fact that they contain new insights. In a slogan, not all systems that we may count as implementations are ‘mere’—informative implementations can point us to new insights into previously postulated structures (Smolensky,, 1988; Smolensky and Legendre,, 2011). Though interpretive work evaluating the compositional models reviewed here is still in its initial stages, it is a promising line of future investigation (Andreas,, 2019; Lepori et al.,, 2023; Lewis et al.,, 2023; McCoy et al.,, 2019; Russin et al.,, 2021; Soulos et al.,, 2020; Todd et al.,, 2024). From a philosophical perspective, further progress on deeper issues about the nature of explanation and implementation will put us in a position to use these findings to resolve the mere implementation objection once and for all.
@@ 4.4 Conclusion
In this article, we have traced the notion of compositionality from its emergence in early analytic philosophy to its present status as a key explanatory construct in contemporary cognitive science. We have emphasized the role that compositionality plays in debates over the explanatory adequacy of neural network models of human cognition and reviewed a range of recent results suggesting that current deep neural networks replicate key behavioral signatures that have motivated the ascription of internal compositional systems to human beings. In particular, these recent results suggest that metalearning (or large-scale pretraining) is a viable approach to reproducing these behavioral signatures within strictly neural architectures. When taken seriously as an approach to modeling human compositionality, the metalearning framework generates concrete predictions about its underlying neural mechanisms and development. DNNs’ recent success also reopens key philosophical questions about the theoretical significance and proper interpretation of neural networks—whether they should be understood as merely implementing classical compositional mechanisms, or as sources of novel insight that may motivate revisions to classical theories. These pressing questions take us beyond what neural networks can do, impelling us to ask what they mean for our understanding of the human mind. We expect that resolving them will require an interdisciplinary approach and increased philosophical engagement.
@@ 4.5 Acknowledgments
We would like to thank Lotem Elber-Dorozko, Richard Kimberley Heck, Christopher Hill, Elizabeth Miller, Joshua Schechter, Roman Feiman, Michael J. Frank, Ellie Pavlick, Randall O’Reilly, Edouard Machery, Suraj Anand, Steven Frankland, Taylor Webb, Stavros Orfeas Zormpalas, Felipe De Brigard, Walter Sinnott-Armstrong, and the participants in the Brown University Dissertation Workshop, the Philosophy and Cognitive Science of Deep Learning reading group, and the 2024 SSNAP Workshop for helpful comments and discussions. JR was supported by NIH NIGMS COBRE grant GR5271961.